OpenStudy (anonymous):

Find the missing exponent. Set up an equation and solve. : (3a^3)^n · 2a^3 = 18a^9 N = missing exponent. Please explain so I can understand.

OpenStudy (kinggeorge):

The original statement is\[[3(a^n)]^n \cdot 2(a^3)=18(a^9)\]correct?

OpenStudy (kinggeorge):

with a 3 instead of an n for that very first \(a^n\)

OpenStudy (anonymous):

No, I meant: \[(3a ^{3})^{n} · 2a ^{3} = 18a ^{9}\]

OpenStudy (kinggeorge):

Then\[(3a^3)^n = (3^n)((a^3)^n)= 3^n \cdot a^{3n}\]If we multiply that by \(2(a^3)\), we get\[ 3^n \cdot a^{3n}\ \cdot 2(a^3)=2\cdot 3^n \cdot a^{3n+3}\]Finally, we set this equal to \(18a^9\) so that \[2\cdot 3^n \cdot a^{3n+3}=18a^9\]Here, notice that \(3n+3=9\)since the exponents on \(a\) must be the same. Solving for \(n\), we get that \(n=2\).If we check our answer, \[2\cdot 3^2 \cdot a^{3\cdot2+3}=2\cdot9\cdot a^9=18a^9\]Since this works, we are done, and \(n=2\).

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (kinggeorge):

You're welcome.